OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? 2x^2 + 3x = -1 a) 0 b) 1 c) 17

OpenStudy (amistre64):

2x^2 +3x = -1 2x^2 +3x +1 = 0 b=3; a=2; and c=1; (3)^2-4(2)(1) = 9-8 = 1

OpenStudy (anonymous):

2x^2 + 3x + 1 = 0 delta = b^2 - 4ac = 9 - 8 = 1 x1 = 3+1/2 = 2 x2 = 3-1/2 = 1 S={1,2} am i not right ?!

OpenStudy (radar):

It looks lilke it could be factored: (2x+1) (x+1)=0 Now for that to equal zero, then one or both of the factors must erqual zero. or (2x+1)=0 then x=-1/2 or (x+1)=0 then x=-1 those are the two roots of your equation.

OpenStudy (radar):

Please excuse the typos

OpenStudy (anonymous):

that looks a good solutions :)